Glentel has sales of $2,780,000. The cost of goods sold is equal to 68 percent of sales. The company has an average inventory of $350,000. How many days on average does it take the company to sell its inventory?

51.32 days

98.12 days

87.40 days

65.46 days

79.12 days
Answer #1

COGS =68%*Sales =68%*2780000=1890400
Average Inventory =350000
Average inventory days =365/(COGS/Inventory)=365/(1890400/350000) =67.58
Option d is correct option as it is closest value